OpenStudy (anonymous):

450 is 75.2% of what number

OpenStudy (ffrocks3):

Set up a ratio: 75.2%=752/1000 450/x=752/1000 Cross multiply 752x=450000 x=598.404

OpenStudy (ffrocks3):

Hope That Helps!

OpenStudy (anonymous):

thank you it did